# Monitored long-range interacting systems: spin-wave theory for quantum trajectories

### v1

@@ V.1 From continuous to discrete monitored dynamics
Figure 8: Benchmark of the quantum-classical measurement protocol. We compare the time evolution of the trajectory average of different correlators (see legend), the quantum-classical correlator $\overline{\langle\hat{S^{z}}\rangle_{Q}\langle\hat{S^{z}}\rangle_{C}}$, the quantum-quantum correlator $\overline{\langle\hat{S^{z}}\rangle_{Q}\langle\hat{S^{z}}\rangle_{Q}}$ and the classical-classical correlator $\overline{\langle\hat{S^{z}}\rangle_{C}\langle\hat{S^{z}}\rangle_{C}}$. Here, trajectory-wise expectation labeled $C$ (classical) is the spin-wave prediction, and the $Q$ (quantum) result is given by the exact simulation with the same realization of the binary noise. The brute-force simulation of a final projective measurement at $\kappa t=60$ is also shown (marked with a cross), where the measured results of the quantum-classical cross-correlated probe $o_{\bm{m}}\langle\hat{S}^{z}_{\bm{m}}\rangle_{C}$ are averaged over 480 runs of the experiment (with different noise realizations). The different panels correspond to different total spin numbers $S$ (see annotation). Parameters: $\omega=1.3\kappa$, $J=0.1\kappa$.
V.2 Quantum-classical observables
To overcome the large post-selection overhead, we combine classical simulation with the output from a quantum simulation. We consider an experiment with $\mathcal{M}\gg 1$ (weak) measurement steps and set the measurement outcome history ${\bm{m}}$. As previously elaborated, the probability of obtaining ${\bm{m}}$ twice is exponentially small in system size and time scale. Suppose we perform a final projective measurement of an operator $\hat{O}$ over $\hat{\rho}({{\bm{m}}})$, i.e. the state conditioned on the measurement history ${\bm{m}}$. This measurement is disruptive and will collapse $\hat{\rho}$ onto the eigenspace of the measurement outcome $o_{{\bm{m}}}$. For example, in the case where the observable is chosen to be $\hat{O}=\hat{S}^{z}$, the single-shot measurement outcome $o_{\bm{m}}\in\{-S,\cdots,S\}$ is one of the $12S+$ 1 eigenvalues of operator $\hat{S}^{z}$. Averaging over this value recasts the Lindblad prediction, namely $\overline{o_{{\bm{m}}}}={\rm{Tr}}[\hat{\rho}\hat{O}]$ where $\hat{\rho}=\overline{\hat{\rho}({\bm{m}})}$ is the average state over all possible trajectories ${\bm{m}}$. A correlation that is non-linear in the state can be obtained using classical simulations [87]. Fixing the measurement history and position ${\bm{m}}$ in the classical simulation, we obtain the (classical) estimate of the trajectory-wise expectation value $\langle\hat{O}_{{\bm{m}}}\rangle_{C}\equiv{\rm{Tr}}[\hat{\rho}_{C}({\bm{m}})\hat{O}]$ with the label $\bullet_{C}$ denoting the classically computed quantity. For instance, this is obtained in our setup using the solution of Eq. (1) within the spin-wave approximations. We can then cross-correlate the measurement outcome with the classical simulation result to construct the quantum-classical object $o_{\bm{m}}\langle\hat{O}_{{\bm{m}}}\rangle_{C}$, which, when averaged over ${\bm{m}}$ (i.e. over measurement shots),
gives the quantum-classical correlator, where the label $\bullet_{Q}$ denotes the quantity evaluated on the actual quantum state in the experiment. In the ideal case where the classical simulation is exact, the quantum and classical quantities should coincide and the correlator on the right-hand side of Eq. (55) is a nonlinear function of the state conditioned on the measurement history. This was recently shown to reproduce the measurement-induced transition in a variety of experiments [74], provided simulating $\langle\hat{O}_{\bm{m}}\rangle_{C}$ is easy. The SWQT framework therefore is a compelling toolbox to investigate quantum-classical correlations. In particular, when the semi-classical approximation holds, we expect the data to be reliable with errors $O(1/S)$.
Figure 9: Steady-state trajectory-averaged correlator $\overline{\langle\hat{S}^{z}\rangle\langle\hat{S}^{z}\rangle}$ as a function of the drive $\omega$ for different total spin numbers $S$ (see legend). Panel (a)-(c) show the results for the the quantum-classical correlator $\overline{\langle\hat{S^{z}}\rangle_{Q}\langle\hat{S^{z}}\rangle_{C}}$, the quantum-quantum correlator $\overline{\langle\hat{S^{z}}\rangle_{Q}\langle\hat{S^{z}}\rangle_{Q}}$ and the classical-classical correlator $\overline{\langle\hat{S^{z}}\rangle_{C}\langle\hat{S^{z}}\rangle_{C}}$ respectively. The steady-state value is computed as the average from $\kappa t=50$ to $\kappa t=100$, i.e. over the tail of the dynamics.
To benchmark this quantum-classical measurement protocol, let us consider again the infinite-range case ($\alpha=0$) of the spin model introduced in Sec. II.1, such that numerically exact integration of the stochastic master equation is affordable for sufficiently large $S$ in order for the spin-wave approximations to yield valid results. We refer to the quantities obtained with the exact solution as the “quantum” ones, to simulate what one would get from an actual experiment. The “classical” quantities are then those given by the spin-wave solution using the same noise realization ${\bm{m}}$ as in the corresponding exact simulation (“quantum”) run. As justified in Sec. V.1, we adopt the binarized noise in both the exact and the spin-wave simulations, in order to mimic an experiment on discrete-time digital quantum simulators where the weak monitoring is implemented with the ancilla qubits. we fix the interaction strength to $J=0.1\kappa$, which is the value considered in Sec. IV, and choose the observable for the final projective measurement to be $\hat{O}=\hat{S}^{z}$. The estimator of the statistical average $\overline{o_{\bm{m}}\langle\hat{S}^{z}_{\bm{m}}\rangle_{C}}$ is then the quantum-classical correlator $\overline{\langle\hat{S}^{z}\rangle_{Q}\langle\hat{S}^{z}\rangle_{C}}$ (we omit the subscript ${\bm{m}}$ for a less heavy notation). In the ideal scenario where the classical simulation is also exact as the quantum one along any trajectory, this correlator should be equal to both the classical-classical correlator $\overline{\langle\hat{S}^{z}\rangle_{C}\langle\hat{S}^{z}\rangle_{C}}$ and the quantum-quantum correlator $\overline{\langle\hat{S}^{z}\rangle_{Q}\langle\hat{S}^{z}\rangle_{Q}}$. In Fig. 8, we show an example of the time evolution of these three correlators at a given driving amplitude $\omega=1.3\kappa$ for different system sizes (indicated by the total spin number $S$). Due to the imperfection of the spin-wave approximations along single trajectories, the quantum-classical correlator gradually deviates from the quantum-quantum and the classical-classical ones, yet still captures qualitatively the dynamics. We also show the brute force simulation result of a projective measurement at a “final” time $\kappa t=60$, where a number of 480 shots of measurement outcomes $o_{\bm{m}}$ are collected and cross-correlated with the classical simulation result to construct the quantum-classical object $\overline{o_{\bm{m}}\langle\hat{S}^{z}_{\bm{m}}\rangle_{C}}$, which agrees well with its estimator $\overline{\langle\hat{S}^{z}\rangle_{Q}\langle\hat{S}^{z}\rangle_{C}}$. In Fig. 9, we show the steady-state behavior of the three correlators discussed above, as a function of the drive $\omega$ for different system sizes. Importantly, the quantum-classical correlation $\overline{\langle\hat{S}^{z}\rangle_{Q}\langle\hat{S}^{z}\rangle_{C}}$ successfully captures the signature of the phase transition. This demonstrates that the efficient classical simulation enabled by our spin-wave framework provides access to quantum-classical observables, which can probe nonlinear properties of the state without the post-selection issue.
VI Conclusions and outlook
In this work, we have proposed a stochastic spin-wave theory along quantum trajectories of monitored long-range spin systems, which serves both as an algorithm for solving the Lindblad master equation and also as a method for investigating monitored dynamics on the level of trajectories with direct access to any quantity of the monitored state, such as entanglement, within the spin-wave approximations. Our technique has been carefully benchmarked in regimes where numerically exact solutions are possible, showing good accuracy even on the level of single trajectories. We further demonstrated the method on a large number of spins with power-law interactions, a regime well beyond the reach of direct simulations, and studied the entanglement dynamics in different regimes of the interaction range. Given the generality of our formalism, we expect the proposed framework of spin-wave quantum trajectories to pave the way for the investigation of a large class of long-range interacting systems.
Despite the demonstrated performance of our method, there are several interesting directions for future studies. For example, one can consider terms in the Holstein-Primakoff expansion beyond the leading order to include more nonlinear effects, which should extend the validity of the theory in the regime of relatively high spin-wave densities. We also expect better overall single-trajectory accuracy with the inclusion of higher-order spin-wave corrections, which is essential in improving the performance of the quantum-classical measurement protocol in avoiding the post-selection problem. On the other hand, different unravellings of the Lindblad master equation can also be explored, including the relevant case of quantum jumps modeling the coupling to photodetectors.
The long-range spin model we adopted for the demonstration of the spin-wave theory can be readily realized in current experimental platforms [121, 23], and we expect the entanglement phase transition to be readily observable in experiments without the post-selection problem. The reason is two-fold. Firstly, as we demonstrated with the spin-wave simulations, the entanglement dynamics features fast saturation times when the interaction is sufficiently long-range, which, together with the collective nature of the dissipation, allows the brute-force post-selection of quantum trajectories with a mitigated, non-exponential, overhead. Finally, in a much broader context, when the semi-classical approximations are valid – including the relevant case of sufficiently long-range interactions, the efficient classical simulation of spin-wave quantum trajectories allows to probe nonlinear quantities of the state via quantum-classical cross-correlated observables, which does not suffer from the post-selection problem, thus opening up promising avenues in the experimental detection of monitored phases in long-range systems.
@@ Acknowledgments
Appendix A Proof of the unraveling of the nondiagonal Lindblad master equation
The positive semidefinite matrix $f_{ij}$ in the nondiagonal master equation (6) can be diagonalized with a unitary transformation $u$:
$f=u\kappa u^{\dagger}\,,$ (56)
@@ Appendix A Proof of the unraveling of the nondiagonal Lindblad master equation
$\kappa\equiv\mathrm{diag}(\kappa_{1},\ldots,\kappa_{N})\,.$ (57)
@@
the Liouvillian (6) becomes
$\mathcal{L}(\hat{\rho})=-{\rm{i}}[\hat{H},\hat{\rho}]+\sum_{i}\kappa_{i}\mathcal{D}[\hat{A}_{i}](\hat{\rho})\,,$ (59)
@@
$\mathcal{D}[\hat{A}_{i}](\hat{\rho})\equiv\hat{A}_{i}\hat{\rho}\hat{A}_{i}^{\dagger}-\dfrac{1}{2}\mathopen{}\mathclose{{}\left\{\hat{A}_{i}^{\dagger}\hat{A}_{i},\hat{\rho}}\right\}$ (60)
is the standard (diagonal) dissipator. The unraveling (1) becomes
${\rm{d}}\hat{\rho}$ $={\rm{d}}t\mathcal{L}(\hat{\rho})$ (61)
@@
which implies that each ${\rm{d}}Z_{i}$ is a normalized complex Wiener process and independent from each other. Eq. (61), being equivalent to Eq. (1), is the standard (diagonal-form) quantum state diffusion unraveling describing a heterodyne detection process, where the operators $\hat{A}_{i}$ are being continuously monitored at rates $\kappa_{i}$ respectively.
@@
Appendix B Mean-field equations for the power-law spin model
We derive in this appendix the mean-field equations of the power-law spin model defined in Sec. II.1. We denote the average magnetization with
$m_{\mu}\equiv\dfrac{\langle\hat{S}^{\mu}\rangle}{S}=\dfrac{1}{N}\sum_{i=1}^{N}\langle\hat{\sigma}^{\mu}_{i}\rangle\,,\quad\mu\in\{x,y,z\}\,,$ (64)
@@ Appendix B Mean-field equations for the power-law spin model
$\dfrac{{\rm{d}}\langle\hat{O}\rangle}{{\rm{d}}t}={}$ ${\rm{i}}\mathopen{}\mathclose{{}\left\langle[\hat{H},\hat{O}]}\right\rangle+\dfrac{\kappa}{S}\mathopen{}\mathclose{{}\left\langle\hat{S}^{+}\hat{O}\hat{S}^{-}-\dfrac{1}{2}\mathopen{}\mathclose{{}\left\{\hat{S}^{+}\hat{S}^{-},\hat{O}}\right\}}\right\rangle\,.$ (65)
@@
$m_{z}=-\sqrt{1-\dfrac{\omega^{2}}{16J^{2}+\kappa^{2}}}\,,$ (67)
giving Eq. (15) in the main text. Beyond the critical point $\omega_{\mathrm{MF}}^{(c)}=\sqrt{16J^{2}+\kappa^{2}}$, this stationary solution no longer exists and the magnetization admits only permanently oscillating solutions, which corresponds to a time-crystal phase.
Appendix C Expression for the entanglement entropy under the Gaussian approximation
@@
$\Xi_{ij}\equiv\frac{1}{2}\langle\hat{r}_{i}\hat{r}_{j}+\hat{r}_{j}\hat{r}_{i}\rangle-\langle\hat{r}_{i}\rangle\langle\hat{r}_{j}\rangle\,,$ (68)
@@
$S_{E}$ $=\sum_{i}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\mu_{i}+\dfrac{1}{2}}\right)\log\mathopen{}\mathclose{{}\left(\mu_{i}+\dfrac{1}{2}}\right)}\right.$ (69)
@@
$\mathbf{\Omega}=\begin{pmatrix}\mathbf{0}&\mathbf{I}_{M}\\ -\mathbf{I}_{M}&\mathbf{0}\end{pmatrix}\,,$ (70)
@@
Figure 10: Results for the power-law interacting Tavis-Cummings model with $\alpha=0.1$, $J=0.01\kappa$ and $\lambda=0.2\kappa$ of different system sizes (colorbar shared across all panels). (a) Steady-state average $z$ magnetization as a function of the drive $\omega$. The vertical dotted line marks the critical point predicted by the mean-field theory $\omega^{(c)}_{\mathrm{MF}}\simeq 0.089\kappa$. (b) Steady-state of the trajectory-averaged entanglement entropy (between the spins and the cavity) as a function of $\omega$. (c) Steady-state spin-wave density. (d) Dynamics of the trajectory-averaged spin-cavity entanglement entropy for a driving value $\omega=0.09\kappa$, in linear-log scale.
@@
Figure 11: Comparison between solutions given by the two different spin-wave methods. (a) Results from the spin-wave quantum trajectories (SWQT) on the expectation value of the collective spin vector in the time-crystal regime with parameters $\omega_{F}=1.25\kappa$, $J=0.1\kappa$ and $S=64$ (the same as Fig. 3). (b) The solution given by the deterministic spin-wave theory (DSW) with the same parameters. (c) Spin-wave densities associated with the two solutions.
Appendix D Generalization to spin-boson systems

### v2

@@ IV.2 From continuous to discrete monitored dynamics
Data availability
The data generated in this study has been deposited in the Zenodo public folder [74].
Code availability
The computer code developed in this study is available from the authors upon reasonable request. A demo is also provided in the Code Ocean capsule [75].
@@ Acknowledgments
Author contributions
Z.L., A.D., X.T. and R.F. contributed to the conception and the implementation of the research, to the discussion of the results, and to the writing of the manuscript.
Competing interests
The authors declare no competing interests.
Supplementary Information for Monitored long-range interacting systems: spin-wave theory for quantum trajectories
Supplementary Note A Proof of the unraveling of the nondiagonal Lindblad master equation
In the nondiagonal master equation (6) of the main text, the positive semidefinite matrix $f_{ij}$ can be diagonalized with a unitary transformation $u$:
$f=u\kappa u^{\dagger}\,,$ (S1)
@@
$\kappa\equiv\mathrm{diag}(\kappa_{1},\ldots,\kappa_{N})\,.$ (S2)
@@
the Liouvillian (6) in the main text becomes
$\mathcal{L}(\hat{\rho})=-{\rm{i}}[\hat{H},\hat{\rho}]+\sum_{i}\kappa_{i}\mathcal{D}[\hat{A}_{i}](\hat{\rho})\,,$ (S4)
@@
$\mathcal{D}[\hat{A}_{i}](\hat{\rho})\equiv\hat{A}_{i}\hat{\rho}\hat{A}_{i}^{\dagger}-\dfrac{1}{2}\mathopen{}\mathclose{{}\left\{\hat{A}_{i}^{\dagger}\hat{A}_{i},\hat{\rho}}\right\}$ (S5)
is the standard (diagonal) dissipator. The unraveling (1) in the main text becomes
${\rm{d}}\hat{\rho}$ $={\rm{d}}t\mathcal{L}(\hat{\rho})$ (S6)
@@
which implies that each ${\rm{d}}Z_{i}$ is a normalized complex Wiener process and independent from each other. Eq. (S6), being equivalent to Eq. (1) of the main text, is the standard (diagonal-form) quantum state diffusion unraveling describing a heterodyne detection process, where the operators $\hat{A}_{i}$ are being continuously monitored at rates $\kappa_{i}$ respectively.
@@
Supplementary Note B Mean-field equations for the power-law spin model
We derive in this Section the mean-field equations of the power-law spin model defined in Sec. II.1.1 of the main text. We denote the average magnetization with
$m_{\mu}\equiv\dfrac{\langle\hat{S}^{\mu}\rangle}{S}=\dfrac{1}{N}\sum_{i=1}^{N}\langle\hat{\sigma}^{\mu}_{i}\rangle\,,\quad\mu\in\{x,y,z\}\,,$ (S9)
@@
$\dfrac{{\rm{d}}\langle\hat{O}\rangle}{{\rm{d}}t}={}$ ${\rm{i}}\mathopen{}\mathclose{{}\left\langle[\hat{H},\hat{O}]}\right\rangle+\dfrac{\kappa}{S}\mathopen{}\mathclose{{}\left\langle\hat{S}^{+}\hat{O}\hat{S}^{-}-\dfrac{1}{2}\mathopen{}\mathclose{{}\left\{\hat{S}^{+}\hat{S}^{-},\hat{O}}\right\}}\right\rangle\,.$ (S10)
@@
$m_{z}=-\sqrt{1-\dfrac{\omega^{2}}{16J^{2}+\kappa^{2}}}\,.$ (S12)
Beyond the critical point $\omega_{\mathrm{MF}}^{(c)}=\sqrt{16J^{2}+\kappa^{2}}$, this stationary solution no longer exists and the magnetization admits only permanently oscillating solutions with zero mean when averaged over long times. The mean-field theory therefore predicts a continuous phase transition from a stationary phase to a time-crystal phase on the level of the Lindblad dynamics of the average state.
Supplementary Note C Expression for the entanglement entropy under the Gaussian approximation
@@
$\Xi_{ij}\equiv\frac{1}{2}\langle\hat{r}_{i}\hat{r}_{j}+\hat{r}_{j}\hat{r}_{i}\rangle-\langle\hat{r}_{i}\rangle\langle\hat{r}_{j}\rangle\,,$ (S13)
@@
$S_{E}$ $=\sum_{i}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\mu_{i}+\dfrac{1}{2}}\right)\log\mathopen{}\mathclose{{}\left(\mu_{i}+\dfrac{1}{2}}\right)}\right.$ (S14)
@@
$\mathbf{\Omega}=\begin{pmatrix}\mathbf{0}&\mathbf{I}_{M}\\ -\mathbf{I}_{M}&\mathbf{0}\end{pmatrix}\,,$ (S15)
@@
Supplementary Note D Additional numerical results
D.1 Benchmark of SWQT on the steady-state of the infinite-range model
Figure S1: Benchmark of the spin-wave method on the steady state of the collective model with $S=64$ and $J=0.1\kappa$. (a) Expectation of the collective $z$-magnetization as a function of the drive $\omega$ obtained with the spin-wave quantum trajectories and the exact solution (see legend). The mean-field solution (see supplementary) is marked with the dashed line. (b) Trajectory-averaged steady-state half-chain entanglement entropy obtained with the two methods [see legend in (a)]. (c) Trajectory-averaged spin-wave density $\overline{\epsilon}$ in the steady state.
In Fig. S1, we compare the steady-state of trajectory average quantities from the spin-wave solutions on the infinite-range model as considered in Sec. II.3.1 of the main text, against the exact ones, for a wide range of the drive $\omega$. Panel (a) shows the magnetization and panel (b) shows the half-system entanglement entropy. Panel (c) shows the spin-wave density $\overline{\epsilon}$ associated with the spin-wave solutions. The peak in $\overline{\epsilon}$ corresponds to the maximum in the entanglement entropy [in panel (b)], where the difference between the spin-wave and exact solutions is also more pronounced. This illustrates the significance of the control parameter $\overline{\epsilon}$ as a signal for the validity of the spin-wave theory. On the other hand, the qualitative behavior of the entanglement is correctly captured despite the relatively high spin-wave density at its peak, and the magnetization predicted by the spin-wave method [panel (a)] remains accurate for all values of $\omega$ considered.
D.2 Comparison with other existing spin-wave methods
@@
Figure S2: Comparison between solutions given by the two different spin-wave methods. (a) Results from the spin-wave quantum trajectories (SWQT) on the expectation value of the collective spin vector in the time-crystal regime with parameters $\omega_{F}=1.25\kappa$, $J=0.1\kappa$ and $S=64$. (b) The solution given by the deterministic spin-wave theory (DSW) with the same parameters. (c) Spin-wave densities associated with the two solutions.
In this section, we provide a comparison between our method of spin-wave quantum trajectories (SWQT) and the method proposed in [58], which is a deterministic spin-wave method for dissipative systems (referred to as “DSW” in this section). The latter assumes the same approximations as in our method (i.e. lowest-order Holstein-Primakoff expansion and Gaussian approximation) except that their approximations are performed on the level of the averaged state (i.e. the density matrix) and their dynamical evolution is derived from the deterministic Lindblad master equation.
For illustration purposes, let us consider again the collective spin model defined in Sec. II.3.1 of the main text. The solutions given by the two methods are shown in Fig. S2 for comparison. Panel (a) shows the SWQT results for the parameters $\omega_{F}=1.25\kappa$, $J=0.1\kappa$, $S=64$. The solution with the same parameters given by DSW is shown in panel (b), which is by no means close to the exact solution. As explained in the main text, this is a result of the highly mixed nature of the average state that the spin-wave approximations fail to capture. Panel (c) shows the spin-wave densities associated with the two solutions, where the SWQT has a spin-wave density that is smaller by several orders of magnitude, which is consistent with its accuracy compared to the exact solution.
D.3 Additional results for the power-law spin model with $\alpha=0.2$ in 1D
Figure S3: Results for the power-law interacting model in the case of $\alpha=0.2$, $J=0.1\kappa$ of different system sizes in 1D (colorbar shared across panels). (a) The scaling of the steady-state average spin-wave density $\overline{\epsilon}$ as a function of the drive $\omega$, in log-linear scale. The vertical dotted lines mark the critical point predicted by the mean-field theory $\omega^{(c)}_{\mathrm{MF}}\simeq 1.077\kappa$. (b) Time evolution of the trajectory-average entanglement entropy at a driving value close to criticality $\omega=1.06\kappa$, in linear-log scale.
Figure S4: Results for the power-law interacting model in the case of $\alpha=1.0$, $J=0.1\kappa$ for different system sizes in 1D (colorbar shared across all panels). (a) Steady-state average $z$ magnetization as a function of the drive $\omega$. The dashed line marks the mean-field solution. (b) Steady-state of the trajectory-averaged half-chain entanglement entropy as a function of $\omega$. The vertical dotted lines mark the critical point predicted by the mean-field theory $\omega^{(c)}_{\mathrm{MF}}\simeq 1.077\kappa$. Inset: scaling of the maximum entropy versus $N$ in linear scale. The dotted line is a linear fit for $N$ ranging from 32 to 128 and we report the coefficient of determination to be $R^{2}=0.9980$. (c) Steady-state spin-wave density in log-linear scale. (d) Dynamics of the trajectory-averaged half-chain entanglement entropy for a driving value $\omega=1.06\kappa$, in linear-log scale.
We study the scaling of the steady-state spin-wave density of the power-law spin model on a 1D chain with $\alpha=0.2$ as considered in Sec. II.3.2 of the main text, as shown in Fig. S3 (a). When the system size $N$ increases, the spin-wave density is suppressed (including its peak value), which is a signature of the long-range nature of the model: in the $N\to\infty$ limit, a vanishing spin-wave density suggests that the system becomes equivalent to a mean-field (infinite-range) one. In Fig. S3 (b), we study the time evolution of the trajectory-averaged entanglement entropy at a driving value $\omega=1.06\kappa$, which is close to where the maximum steady-state entanglement is achieved for the finite system sizes we studied. Note that we are increasing $N$ exponentially between the different considered values, while the entanglement quickly reaches the steady-state value in all cases, suggesting the mitigation of the exponential overhead for observing the entanglement transition via brute-force post selection.
D.4 Results for the power-law spin model with $\alpha=1$ in 1D
To study the effect of the interaction range on the entanglement dynamics in the power-law spin model, we also consider the shorter-range case with $\alpha=1$ on a 1D chain for comparison. The steady-state magnetization $\overline{\langle\hat{S}^{z}\rangle}$, as shown in Fig. S4 (a), exhibits qualitatively similar behavior to the long-range case, while it shows more deviation from the mean-field prediction. This could result from more significant finite-size effects due to short-range interactions. (Note that the mean-field theory does not account for the interaction range parameter $\alpha$ and therefore gives identical predictions regardless of $\alpha$.) In sharp contrast, Fig. S4 (b) shows that the entanglement entropy of the short-range systems grows much faster with the system size $N$. Finite-size scaling suggests a volume law for the maximum entanglement entropy as well as in a vicinity around the critical point. The spin-wave density, as shown in Fig. S4 (c), also presents a qualitatively different behavior from the long-range case. Close to the maximal entanglement, the spin-wave density does not appear to decrease with the system size $N$, and remains around $\overline{\epsilon}\lesssim 0.2$. This suggests that the short-range model does not reduce to a mean-field (infinite-range) one even in the thermodynamic limit of $N\to\infty$, and that higher-order corrections to the theory are expected to have a more significant contribution. Finally, we show in Fig. S4 (d) the time dynamics of the trajectory-averaged entanglement entropy at $\omega=1.06\kappa$. Contrary to the fast saturation of entanglement in the long-range (or infinite-range) regime, the time it takes to reach the steady-state value scales rapidly with $N$. Despite the relatively high spin-wave density at the considered driving, we expect our result to hold qualitatively.
Figure S5: Results for the power-law spin model in 2D with $\alpha=0.2$, $J=0.1\kappa$ for the following trajectory-averaged quantities in the steady state: (a) $z$ magnetization $\overline{\langle\hat{S}^{z}\rangle}/S$ with the mean-field solution marked by the dashed line, (b) half-system entanglement entropy $\overline{S_{E}}$ with its maximum value in the inset, and (c) spin-wave density $\overline{\epsilon}$, as functions of the drive $\omega$ for different system sizes $N=L^{2}$ (see colorbar).
Figure S6: Same as Fig. S5 but for $\alpha=2.0\,$. The dotted line in the inset is a linear fit for $N$ ranging from 36 to 256 and we report the coefficient of determination to be $R^{2}=0.9991$.
Figure S7: Same as Fig. S5 but for $\alpha=3.0\,$. The dotted line in the inset is a linear fit for $N$ ranging from 25 to 100 and we report the coefficient of determination to be $R^{2}=0.9983$.
D.5 Results for the power-law spin model in 2D
In this section, we present additional results on the power-law spin model for a 2D square lattice. Fig. S5 shows the steady-state magnetization [panel (a)], trajectory-averaged half-system entanglement entropy [panel (b)], and the spin-wave density [panel (c)], for the case of $\alpha=0.2$. As the system size $N=L^{2}$ increases, the magnetization shows the (dissipative) phase transition predicted by the mean-field theory. The entanglement also develops a log divergence at the same critical point, accompanied by a maximal spin-wave density (with respect to $\omega$) that decreases with $N$. As the interaction range is sufficiently long, these results are similar to the 1D case with $\alpha=0.2$ discussed in the main text (Sec. II.3.2), and both instances are showing essentially the mean-field (infinite-range) physics. Note that we report results for up to $N=32\times 32=1024$ interacting spins in 2D.
Fig. S6 shows the results for $\alpha=2.0$, which is the boundary case between our identification of long-range and short-range regimes. Similar to the 1D case with $\alpha=1.0$, the entanglement scaling at its maximum becomes asymptotically volume-law. The peak spin-wave density no longer vanishes with $N$, yet still remains bounded and small ($<0.1$), suggesting the validity of the spin-wave approximations. The transition in $\langle\hat{S}^{z}\rangle$ remains qualitatively similar to the mean-field solution (since this transition is driven by the infinite-range dissipation and not the coherent interaction term), with a small deviation originating from the quantum fluctuations close to the transition. These features are more manifest in Fig. S7, where we study a moderately short-range case with $\alpha=3.0\,$. The entanglement exhibits a volume-law scaling again and the magnetization shows more deviation from the mean-field solution due to the increased quantum fluctuations. Contrary to the previous cases, the spin-wave density increases with the system size $N$ for drive values close to and beyond the mean-field critical point. We therefore limit the system sizes studied (up to $N=10\times 10=100$ spins in 2D) such that the spin-wave density remains small around $\overline{\epsilon}\lesssim 0.1$.
Finally, we show in Fig. S8 the drive value at which maximum entanglement is achieved, i.e. $\mathrm{argmax}_{\omega}\{\overline{S_{E}}\}$, for the different cases we considered in both 1D and 2D. As the system size $N$ increases, this drive value becomes asymptotically independent of $N$ and converges to a value close to the mean-field critical point.
Figure S8: The drive that maximizes the steady-state trajectory-averaged entanglement entropy, $\mathrm{argmax}_{\omega}\{\overline{S_{E}}\}$, as a function of the system size $N$ for different interaction ranges $\alpha$. The errorbar is set at the increment at which we scan $\omega$, i.e. $\Delta\omega=0.01\kappa$ in our settings. The dotted horizontal line marks the mean-field critical drive value $\omega_{\mathrm{MF}}^{(c)}$.
D.6 Benchmark of the binary approximation of the Gaussian noise
In this section, we verify the validity of the binary approximation of the Gaussian noise presented in the Methods section by computing the trajectory-averaged quantities including both linear and non-linear functions of the state, and by comparing the results with the exact solution obtained with the Gaussian noise. An example of the benchmark is shown in Fig. S9, where we computed the magnetization $\overline{\langle\hat{S}^{z}\rangle}$, the half-system entanglement entropy $\overline{S_{E}}$ and the nonlinear quantity $\overline{\langle\hat{S}^{z}\rangle^{2}}$ from three different methods: 1) Spin-wave quantum trajectories with binarized noise, 2) exact solution of the stochastic master equation with Gaussian noise and 3) exact solution with the binary approximation of the noise. In all cases, we observe excellent agreement among the results obtained in these different ways.
Figure S9: Benchmark of the binary approximation of the Gaussian noise in the heterodyne-detection unraveling. We compare trajectory-averaged quantities obtained from different methods and different types of noise (see legend), namely 1) the spin-wave quantum trajectory (SWQT) method with binarized noise, 2) the exact solution of the stochastic master equation with Gaussian noise and 3) the exact solution but with binary noise replacing the gaussian noise. Panel (a) compares the results on a linear quantity $\overline{\langle\hat{S}^{z}\rangle}$. Panels (b) and (c) show the benchmark on two different nonlinear quantities, the half-system entanglement entropy $\overline{S_{E}}$ and the squared magnetization $\overline{\langle\hat{S}^{z}\rangle^{2}}$ respectively. Parameters: $\omega=1.3\kappa$, $J=0.1\kappa$, $\alpha=0$ and $S=64$.
Supplementary Note E Generalization to spin-boson systems
Figure S10: Results for the power-law interacting Tavis-Cummings model with $\alpha=0.1$, $J=0.01\kappa$ and $\lambda=0.2\kappa$ of different system sizes (colorbar shared across all panels). (a) Steady-state average $z$ magnetization as a function of the drive $\omega$. The vertical dotted line marks the critical point predicted by the mean-field theory $\omega^{(c)}_{\mathrm{MF}}\simeq 0.089\kappa$. (b) Steady-state of the trajectory-averaged entanglement entropy (between the spins and the cavity) as a function of $\omega$. (c) Steady-state spin-wave density. (d) Dynamics of the trajectory-averaged spin-cavity entanglement entropy for a driving value $\omega=0.09\kappa$, in linear-log scale.